# Liwen Wang1,2,  Yuanyuan Yuan2,  Ao Sun2,  Zongjie Li2,  Pingchuan Ma2,  Daoyuan Wu2,  Shuai Wang2

### v2

@@ Abstract
Visual deep learning (VDL) systems have demonstrated their capability to understand complex image semantics, paving the way for significant real-world applications like image recognition, object detection, and autonomous driving. To evaluate the reliability of VDL, a mainstream approach is software testing, which requires diverse and controllable mutations over image semantics. The rapid development of multi-modal large language models (MLLMs) has introduced revolutionary image mutation potentials through instruction-driven methods. Users can now freely describe desired mutations and let MLLMs generate the mutated images. Hence, parallel to large language models’ (LLMs) recent success in traditional software fuzzing, one may also expect MLLMs to be promising for VDL testing in terms of offering diverse and complex image mutations.
However, the quality of MLLM-produced test inputs in VDL testing remains largely unexplored. We present the first study, aiming to assess MLLMs’ adequacy from 1) the semantic validity of MLLM mutated images, 2) the alignment of MLLM mutated images with their text instructions (prompts), 3) the faithfulness of how different mutations preserve semantics that are ought to remain unchanged, and 4) the effectiveness of detecting VDL faults. With large-scale human studies and quantitative evaluations, we identify MLLM’s promising potentials in expanding the covered semantics of image mutations. Notably, while SoTA MLLMs (e.g., GPT-4V) fail to support or perform worse in editing existing semantics in images (as in traditional mutations like rotation), they generate high-quality test inputs using “semantic-additive” mutations (e.g., “dress a dog with clothes”), which bring extra semantics to images; these were infeasible for past approaches. Hence, we view MLLM-based mutations as a vital complement to traditional mutations, and advocate future VDL testing tasks to combine MLLM-based methods and traditional image mutations for comprehensive and reliable testing.
@@ 1 Introduction
Visual deep learning (VDL) systems comprehend semantics of image contents to enable a wide range of real-world applications (e.g., autonomous driving, medical imaging diagnosis) that were previously human-reliant. However, since the decision rules of VDL are obscure, such as how a VDL system recognizes cars, VDL systems can experience drastic failures in their outputs, which jeopardizes the reliability of VDL systems in safety-critical applications. Software testing has been very effective in detecting VDL failures; it mutates input images and defines testing oracles over varying mutation schemes (e.g., VDL outputs should be invariant to the rotation of the input). A VDL failure can be detected as a violation of the oracle. In recent years, our community has proposed a variety of mutation schemes for VDL testing under different scenarios, such as image classification [62], object detection [57], image captioning [64], and autonomous driving [46, 56, 67].
@@
Inputs of VDL systems are images whose contents have rich semantics. Thus, to evaluate the reliability of VDL systems via software testing, comprehensively covering different image semantics is demanding for the input mutation. Nevertheless, mutating image semantics is fundamentally challenging because the relation between pixel values and most image semantics is unclear and hard to model [65, 72, 52, 27, 42, 37]. Existing methods operating on pixels can only mutate some basic semantics such as brightness and contrast [46]. Also, the mutation must be controllable; an arbitrary mutation itself can violate the testing oracle and induce an invalid testing input. Previous works achieve controllable mutations for more complex semantics using representation learning techniques [71, 67, 52], which require image samples containing the expected mutations and are limited to certain VDL inputs such as driving scenes [67] or face photos [52].
@@
The recent multi-modal large language models (MLLMs) [^1] Since MLLMs are often jointly used with other large models such as Diffusion model [30], CLIP [48], DALLE [3], etc., to minimize confusion, we use MLLMs as a general term for them. We use their own names when referring to their special usage. (e.g., GPT-4V [5]) have liberated the freedom of diverse and controllable mutations for image semantics: users can freely describe the anticipated mutation in natural language and let MLLMs generate the mutated images. This smooth translation between text descriptions and images is shedding light on a new paradigm of image mutation, namely instruction-driven mutation, that supports various traditional mutations (see the introduction in Sec. 3) in a unified prompt-driven manner, and also offers a wide range of new mutations that are hardly achievable by traditional methods. Thus, one may reasonably anticipate that MLLMs can revolutionize VDL testing, in parallel to large language model (LLM)’s emerging adoption and success in our community to boost traditional software testing [15, 60, 55], program repairing [47, 61, 20], and program synthesis [6, 34, 59].
@@
Despite the optimism towards MLLM-empowered VDL testing, the internals of MLLMs (i.e., how they generate outputs) are hard to interpret. Unlike traditional mutations that form constraints (e.g., explicit mathematical formulas; see Sec. 3.1) to control the mutated semantics, MLLMs lack guarantees on whether the mutations are achieved as expected. Notably, due to the underlying deep learning techniques, modern VDL systems do not have effective input checkers. Thus, a semantically invalid image can be accepted by a VDL system to produce a valid (but incorrect) output. Some recent reports have shown that MLLMs can generate semantically invalid images [18, 10, 21], which makes the follow-up testing and the detected VDL failures meaningless (as those failures can hardly occur in real-world scenarios). Thus, it is demanding to systematically evaluate MLLM’s capability in mutating VDL test inputs.
This paper first presents a taxonomy on image semantics at different granularities and reviews how they are covered by traditional mutation methods and the instruction-driven mutation of MLLMs. Then, for each category of semantics, we assess the quality of test inputs generated by traditional mutations and MLLMs. Following common practice in VDL testing [56, 67, 62], quality is measured from 1) the validity of semantics in mutated images, 2) the alignment of mutated images with text instructions, 3) the faithfulness of how different mutations preserve semantics that should be unchanged, and 4) the effectiveness of detecting VDL faults.
We consider ten representative types of mutations and four recent SoTA MLLMs including GPT-4V. Our evaluated datasets cover general image classification [14], fine-grained dog breed identification [2], face recognition [36], and autonomous driving [13]. We conduct large-scale human studies through Amazon Mechanical Turk [1] by hiring 20 Ph.D. students experienced in VDL systems and software testing. They evaluate 2,300 image mutations, and each mutation is rated by two human participants w.r.t. our criteria. We also conduct quantitative evaluations to measure the quality of test inputs. Our results reveal that MLLM-enabled mutations and traditional mutations are complementary in VDL testing: MLLMs cannot replace traditional mutation methods but bring a new dimension to image mutations. Specifically, while MLLMs can alternatively implement traditional mutations with proper text prompts, the re-implemented mutations exhibit low faithfulness and alignment. For example, none of our evaluated SoTA MLLMs can achieve image rotation. On the other hand, although MLLMs cannot edit existing semantics in images as well as traditional mutations, they perform impressively well on “semantic-additive” mutations that add new semantics to the seed images (e.g., “adding glasses for a dog in an image”), which are uniquely enabled by MLLMs. In sum, this paper makes the following key contributions:
  • This paper conducts the first study on the quality of test inputs generated by MLLM for VDL testing. It also studies the role and positioning of MLLM’s instruction-driven mutation w.r.t. traditional mutation paradigms.
  • We present a systematic review of existing image mutations in VDL testing. We also design large-scale human studies and quantitative evaluations to assess the quality of test inputs from the validity, alignment, faithfulness, and effectiveness perspectives.
@@
  • Our findings better position MLLMs and their enabled instruction-driven mutations in VDL testing. While MLLMs cannot well edit existing semantics in images like traditional mutations, they enable adding new semantics, thus expanding VDL testing’s scope.
@@ 2 Background: VDL Testing
Holistically speaking, a VDL system typically consists of a visual understanding module and a decision module. The visual understanding module takes an image as input and extracts features from its content. The decision module yields outputs (according to the specific VDL application domain) based on the extracted features. From this perspective, image classifier, object detector, image captioner, and auto-driving systems, etc., that were widely tested in our community can all be counted as VDL systems [46, 56, 62, 57, 64].
The vanilla way of evaluating VDL systems is comparing their outputs with human annotations over a set of test inputs (i.e., labeled images in standard test datasets). However, manually annotating test inputs is labor-intensive, impeding automated and large-scale evaluations. Software testing has achieved remarkable success in evaluating the reliability of VDL systems. It uses different input mutation schemes to generate diverse test inputs and forms oracles to enable automatic “annotations” for the test inputs. For example, a rescaled image should have the same output as the original image when fed into an object detector. This way, software testing can identify VDL faults as violations of testing oracles, where human annotations are no longer required.
Test Input Generation. Given that VDL processes semantics of image contents, existing works have proposed various mutations to modify image semantics; we elaborate on them in Sec. 3. Note that different from adversarial perturbations [^2] Such perturbations are leveraged by adversarial attacks to generate malicious inputs (i.e., adversarial examples) to fool a VDL system [25]. that add noise to images or change a few pixels, semantic-level mutations in VDL testing holistically alter images to diversify image semantics, exploring different usage scenarios of VDL systems.
Testing Oracle. Metamorphic testing (MT) stands as one mainstream testing paradigm of VDL testing in recent years [57, 64, 56, 62, 66, 67, 65]. MT checks whether a VDL system’s outputs are aligned with the relation defined in an MT testing oracle, when the VDL is processing an input and its mutated version. Unalignment shows that one of the inputs is ill-processed. Since testing oracles are critical to detect VDL faults, input mutations cannot be arbitrary and should properly preserve the oracle; see examples in Sec. 3.
3 Image Mutations
@@
Image Decomposition: A Motivating Example. To understand how VDL works and how existing input mutations are designed, we first decompose the dog image in Fig. 1 into different semantics. Starting from a lower level, the image can be seen as a matrix of pixels, and the magnitude of pixel values determines the brightness and contrast of the image. Besides, the spatial arrangement of pixels reflects geometrical properties of the image, such as the scale and the rotation angle. A reliable VDL system that recognizes dogs should be resilient to changes of these semantics. At a higher level, the image contains a dog and the background grass; the dog itself has rich semantics such as the action and the orientation. Similarly, the semantics in background may vary with the working scenario of the VDL, e.g., the grass may be replaced by snow in a winter scene. Overall, the dog-recognition VDL should focus on the semantics of the dog and be invariant to irrelevant semantics in the background.
@@ 3.1 Explicit & Mathematical Transformations
Pixel-level. Image brightness and contrast can be mutated by uniformly changing pixel values. In general, mutations belonging to this category can be formed as $\mathbf{x}$ op $c$, where $c$ is a scalar value and $\mathbf{x}$ is a matrix denoting the seed image. op decides the mutated semantics. For example, op is the add operation if the mutation changes the image brightness, that is, to increase the brightness of $\mathbf{x}$, all pixel values in $\mathbf{x}$ can be increased by $c$. Similarly, mutating image contrast can be formulated as $\mathbf{x}\times c$, which enlarges the difference between pixel values in $\mathbf{x}$. Examples are in Fig. 1.
Geometrical-level. As geometrical properties are determined by the spatial arrangement of pixels, existing works implement these mutations using affine transformations, which essentially move pixels to new localizations and preserve the image’s parallelism (i.e., two parallel lines remain parallel in the mutated image). Formally, suppose pixels of the image $\mathbf{x}$ locate in a rectangular coordinate and the centering pixel is the original, the $(i,j)$-th pixel’s localization in the mutated image can be calculated as $[i^{\prime},j^{\prime}]^{\intercal}=\mathbf{A}\times[i,j]^{\intercal}+\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ characterize the geometrical mutation. For example, the values of $(\mathbf{A},\mathbf{b})$ of scaling and translation mutations are shown as $(\mathbf{A}_{S},\mathbf{b}_{S})$ and $(\mathbf{A}_{T},\mathbf{b}_{T})$ in Eq. 1, respectively,
$\mathbf{A}_{R}=\begin{bmatrix}\cos(\theta),&-\sin(\theta)\\ \sin(\theta),&\cos(\theta)\\ \end{bmatrix},\mathbf{b}_{R}=\mathbf{0};\;\mathbf{A}_{T}=\mathbf{0},\mathbf{b}_{T}=\begin{bmatrix}t_{x}\\ t_{y}\\ \end{bmatrix}$ (1)
where $\theta$ specifies the rotating angle. $t_{x}$ and $t_{y}$ indicate the vertical and horizontal translation distances. See examples in Fig. 1.
The aforementioned mutations are preferred in early VDL testing works due to their neat and explicit formulations [46, 56, 62]. These mutations are controllable and faithfully preserve semantics that shouldn’t be changed. The generated images are always valid (if the mutation is not iteratively applied). However, such explicit-formed mutations require knowing the precise relation between semantics and each pixel. Currently, only a few mutations are explicit-formed, largely limiting the diversity of covered image semantics.
@@ 3.2 Implicit & Data Exploration
Some recent researches diversify the mutated semantics with data-driven mutations. These mutations implicitly explore patterns in a set of images and then apply the patterns to mutate a seed image.
@@
Style-level. From a high level, each image can be decomposed as its style and content, and VDL systems should primarily focus on the content rather than style. Thus, existing VDL testing works propose different style-level mutations, whose generated images have distinct styles but presumably retain the content of the seed image. The “style” has been studied from two aspects. First, the style can be viewed as describing how the image is produced, e.g., the seed image in Fig. 1 is taken in reality and its style is “natural”. In that sense, the style can be changed by generating an image’s artistic-stylized variants (e.g., an image having the same dog but painted using a Van Gogh style, as in Fig. 1[65]. Also, some works decompose a weather condition as the style and separate it from the main content. The weather conditions are transferred between images to generate test inputs; see Fig. 1. These weather-related mutations are often used to test VDL-based autonomous driving [67].
Existing style-level mutations are implemented using generative adversarial nets (GANs) and their success highly relies on the cycle-consistency [74]: decomposing and transferring styles are only applicable to images of similar content [73]; see results in Sec. 6.2.
Perceptual-level. The latest VDL testing has been focusing on fine-grained semantics of the objective in images, e.g., as shown in Fig. 1, the dog’s eyes and mouth can be mutated to evaluate whether the VDL accurately recognizes different dogs. Such mutations are often known as perceptual-level mutations [65] and aim to edit the status of existing semantics in images [19, 65]. Perceptual-level mutations employ generative models (e.g., GANs [36]) to obtain a semantic-level latent representation for each image. Instead of operating on the pixel values (as in mutations mentioned in Sec. 3.1), they manipulate the latent representation to achieve the mutation.
To control the mutation and generate valid test inputs, perceptual-level mutations rely on a hypothesis: when projected into the latent space using a well-developed generative model, semantically meaningful images concentrate as clusters and each cluster corresponds to images of similar semantics [40, 65]. For instance, if two clusters $a$ and $b$ in the latent space contain left- and right-oriented dogs, respectively, slightly moving an image’s latent coordinate towards the direction $a\leadsto b$ can mutate the dog’s orientation in the image generated using this new latent coordinate [52]. We refer interested readers to [52, 71] for details of perceptual-level mutations.

### v3

# How Multi-Modal LLMs Reshape Visual Deep Learning Testing? A Comprehensive Study Through the Lens of Image Mutation

@@ Abstract
Visual deep learning (VDL) systems have demonstrated their capability to understand complex image semantics, paving the way for significant real-world applications like image recognition, object detection, and autonomous driving. To evaluate the reliability of VDL, a mainstream approach is software testing, which requires diverse mutations over image semantics. The rapid development of multi-modal large language models (MLLMs) has introduced revolutionary image mutation potentials through instruction-driven methods. Users can now freely describe desired mutations and let MLLMs generate the mutated images. Hence, parallel to large language models’ (LLMs) recent success in traditional software fuzzing, one may also expect MLLMs to be promising for VDL testing in terms of offering unified, diverse, and complex image mutations.
However, the quality and applicability of MLLM-based mutations in VDL testing remain largely unexplored. We present the first study, aiming to assess MLLMs’ adequacy from 1) the semantic validity of MLLMs’ mutated images, 2) the alignment of MLLMs’ mutated images with their text instructions (prompts), and 3) the faithfulness of how different mutations preserve semantics that are ought to remain unchanged. With large-scale human studies and quantitative analysis, we identify MLLM’s promising potential in expanding the covered semantics of image mutations. Notably, while SoTA MLLMs (e.g., GPT-4V) fail to support or perform worse in editing existing semantics in images (as in traditional mutations like rotation), they generate high-quality test inputs using “semantic-replacement” mutations (e.g., “dress a dog with clothes”), which bring extra semantics to images and are effective in triggering VDL faults; these were infeasible for past approaches. Hence, we view MLLM-based mutations as a vital complement to traditional mutations, and advocate future VDL testing tasks to combine MLLM-based and traditional image mutations for comprehensive and reliable testing.
Index Terms:
Deep learning testing, metamorphic testing, multi-modal large language models
@@ 1 Introduction
Visual deep learning (VDL) systems comprehend the semantics of image contents to enable a wide range of real-world applications (e.g., autonomous driving, medical imaging diagnosis) that were previously human-reliant. However, VDL systems can experience drastic failures in their outputs due to their obscure decision rules, jeopardizing the reliability of VDL systems in safety-critical applications. Software testing has been very effective in detecting VDL failures; it mutates input images and defines the accompanying testing oracle and numeric validation metric. With the metric validating the achieved mutation, a VDL failure can be detected as a violation of the oracle. In recent years, our community has proposed a variety of mutation schemes for VDL testing under different scenarios, such as image classification [1], object detection [2], image captioning [3], and autonomous driving [4, 5, 6].
@@
Inputs of VDL systems are images whose contents have rich semantics. Thus, to evaluate the reliability of VDL systems via software testing, comprehensively covering different image semantics is demanding for the input mutation. Nevertheless, mutating image semantics is fundamentally challenging because the relation between pixel values and most image semantics is unclear and hard to model [7, 8, 9, 10, 11, 12]; most existing methods primarily mutate some pixel-level semantics such as brightness and contrast [4]. Some recent works support mutating more complex semantics via representation learning techniques [13, 6, 9], which however require paired images containing the expected mutations and only apply to specific domains such as driving scenes [6] or face photos [9].
@@
The recent multi-modal large language models (MLLMs) [^1] Since MLLMs are often jointly used with other large models such as Diffusion model [14], CLIP [15], DALLE [16], etc., to minimize confusion, we use MLLMs as a general term for them. We use their own names when referring to their special usage., e.g., GPT-4V [17], have liberated the freedom of mutations for image semantics: users can freely describe the anticipated mutation in natural language and let MLLMs generate the mutated images. This new instruction-driven mutation paradigm is shedding light on (1) supporting different traditional mutations in a unified prompt-driven manner, and therefore, one may expect to easily understand, maintain, extend, and disseminate various input mutation methods in natural language texts. (2) Moreover, it also shows promising potential in offering a wide range of new mutations that are hardly achievable by traditional methods, based on the high natural language comprehension ability of LLMs, and high image synthesis capabilities of advanced vision backends (e.g., Diffusion model [14]). Thus, one may reasonably anticipate that MLLMs can revolutionize VDL testing, in parallel to large language model (LLM)’s emerging adoption and success in our community to boost traditional software testing [18, 19, 20], program repairing [21, 22, 23], and program synthesis [24, 25, 26].
@@
Nevertheless, unlike traditional methods that explicitly form the mutations (e.g., via mathematical formulas; see Sec. 3.1), it is inherently challenging to interpret how MLLMs generate mutated outputs. Some recent reports show that MLLMs can generate semantically invalid images [27, 28, 29], which makes the follow-up testing meaningless, as those detected “failures” are induced by the VDL system’s incorrect usages, not its faults. Thus, it is demanding to systematically evaluate MLLM’s capability in mutating images.
Following common practice in VDL testing [5, 6, 1], we first measure the quality of MLLM-generated test inputs from: 1$\rangle$ the validity of the overall semantics in mutated images, 2$\rangle$ the alignment of mutated images with text instructions, and 3$\rangle$ the faithfulness of how different mutations preserve semantics that should be unchanged. This is achieved via a large-scale human evaluation on Amazon Mechanical Turk [30], with 20 Ph.D. students experienced in VDL systems and software testing. Based on the quality assessment, we then study MLLM’s mutation capabilities by answering two research questions (RQs). RQ1: Can MLLMs unify different traditional input mutations? RQ2: What benefits can MLLMs bring to input mutations in VDL testing? Since input validation is crucial to VDL testing, we further study if existing validation metrics apply to MLLM-based mutations (RQ3).
We consider two recently proposed (the only two to our knowledge) pipelines of leveraging MLLMs for image mutations; both of them implement optimizations for prompts and image generation modules to provide out-of-the-box usages. Our evaluations incorporate four SoTA MLLMs (including GPT-4V) into these pipelines to achieve 10 representative mutations, and the evaluated datasets cover general image classification [31], fine-grained dog breed identification [32], face recognition [33], and autonomous driving [34].
Our study reveals that MLLM-enabled mutations and traditional mutations are complementary in VDL testing: MLLMs cannot unify traditional mutations (which primarily edit the status of semantics that are already in the seed images) and exhibit low faithfulness and alignment when implementing them, e.g., none of our evaluated SoTA MLLMs can achieve image rotation. However, MLLMs bring a new dimension to image mutation: they perform impressively well on “semantic-replacement” mutations that replace semantics in the seed images (e.g., “change the background to a library”); these are uniquely enabled by MLLMs. Besides, our results show that the two mutation pipelines of MLLMs may fit distinct scenarios, and call for careful calibrations when using them. Moreover, we find that existing validation metrics are less applicable to MLLM-based mutations due to the diversified mutants and finer-grained operations, and suggest potential improvements. In sum, this paper makes the following key contributions:
  • This paper conducts the first study on the quality of test inputs generated by MLLM for VDL testing, and designs a large-scale human evaluation from the validity, alignment, and faithfulness perspectives.
  • Our results reveal the (in-)capabilities of MLLMs in image mutation, the limitation of existing validation metrics, and the distinct suitability of different mutation pipelines that are designed for MLLMs. We also summarize the lessons learned and provide suggestions for future improvements.
@@
  • Our findings better position MLLMs and their enabled instruction-driven mutations in VDL testing. While MLLMs cannot edit existing semantics in images well like traditional mutations, they enable replacing semantics with new ones, being complementary to traditional mutations and expanding VDL testing’s scope.
@@ 2 Background: VDL Testing
In general, a visual deep learning (VDL) system typically consists of a visual understanding module and a decision module. The visual understanding module takes an image as input and extracts features from its content. The decision module yields outputs (according to the specific VDL application domain) based on the extracted features. From this perspective, image classifiers, object detectors, image captioners, and auto-driving systems, etc., that were widely tested in our community can all be counted as VDL systems [1, 2, 3, 4, 5].
The vanilla way of evaluating VDL systems is comparing their outputs with human annotations over a set of test inputs (i.e., labeled images in standard test datasets). However, manually annotating test inputs is labor-intensive, impeding automated and large-scale evaluations. Software testing has achieved remarkable success in evaluating the reliability of VDL systems. It uses different input mutation schemes to generate diverse test inputs and forms oracles to enable automatic “annotations” for the test inputs. For example, a slightly rescaled image (e.g., no less than $50\%$, as determined by the validation metric) should have the same output as the original image when fed into an object detector. This way, software testing can identify VDL faults as violations of testing oracles, where human annotations are no longer required.
Test Input Generation and Validation. Since VDL processes the semantics of image contents, various mutations and validation metrics have been proposed to generate legitimate test inputs (see details in Sec. 3). In short, the mutations broadly modify different semantics to explore potential application scenarios of VDL systems. The numeric validation metrics quantify the extents of achieved mutations and the validity of test inputs, in order to control the mutation and ensure meaningful testing results. Note that different from adversarial perturbations [^2] Such perturbations are leveraged by adversarial attacks to generate malicious inputs (i.e., adversarial examples) to fool a VDL system [35]. that add invisible noise to images and study the “worst-case” attack surfaces of VDL systems, mutations employed in VDL testing holistically alter images to diversify image semantics, exploring different usage scenarios of VDL systems.
Testing Oracle. Metamorphic testing (MT) stands out as one mainstream testing paradigm of VDL testing in recent years [2, 3, 5, 1, 36, 6, 7]. MT checks whether a VDL system’s outputs are aligned with the relation (e.g., an equality relation) defined in an MT testing oracle, when the VDL is processing an input and its mutated version. Unalignment shows that one of the inputs is ill-processed and a VDL fault is triggered. Since testing oracles are critical to detect VDL faults, input mutations themselves should properly preserve the oracles (e.g., within specific semantic ranges, as determined by the validation metric); see examples in Sec. 3.
3 Image Mutations and Validation
@@
Image Decomposition: A Motivating Example. To understand VDL and the design of input mutations, we first decompose the dog image in Fig. 1 into different semantics. Starting from a lower level, the image is a matrix of pixels, and the magnitude of pixel values determines the brightness and contrast of the image. Besides, the spatial arrangement of pixels reflects geometrical properties of the image, such as the scale and the rotation angle. A reliable VDL system that recognizes dogs should be resilient to changes of these semantics. At a higher level, the image contains a dog and the background grass; the dog itself has rich semantics such as the action and the orientation. Similarly, the semantics in background may vary with the working scenario of the VDL, e.g., the grass may be replaced by snow in a winter scene. Overall, the dog-recognition VDL should focus on the semantics of the dog and be invariant to irrelevant semantics in the background.
@@ 3.1 Explicit & Mathematical Transformations
Pixel-level. Mutations belonging to this category can be formed as uniformly changing pixel values via $\mathbf{x}$ op $c$, where $c$ is a scalar value and $\mathbf{x}$ is a matrix denoting the seed image. op decides the mutated semantics. For example, op is the add operation if the mutation changes the image brightness, i.e., adding all pixel values in $\mathbf{x}$ with a positive number $c$ increases $\mathbf{x}$’s brightness. Similarly, mutating image contrast can be formulated as $\mathbf{x}\times c$, which enlarges the difference between pixel values in $\mathbf{x}$. Representative examples are shown in Fig. 1.
Geometrical-level. As geometrical properties are determined by the spatial arrangement of pixels, existing works implement these mutations using affine transformations, which essentially move pixels to new localizations and preserve the image’s parallelism (i.e., two parallel lines remain parallel in the mutated image). Formally, suppose pixels of the image $\mathbf{x}$ are in a rectangular coordinate and the centering pixel is the origin, the $(i,j)$-th pixel’s localization in the mutated image can be calculated as $[i^{\prime},j^{\prime}]^{\intercal}=\mathbf{A}\times[i,j]^{\intercal}+\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ characterize the geometrical mutation. For example, $(\mathbf{A}_{S},\mathbf{b}_{S})$ and $(\mathbf{A}_{T},\mathbf{b}_{T})$, w.r.t. scaling and translation mutations, are shown in Eq. 1,
$\mathbf{A}_{R}=\begin{bmatrix}\cos\theta,&-\sin\theta\\ \sin\theta,&\cos\theta\\ \end{bmatrix},\mathbf{b}_{R}=\mathbf{0};\;\mathbf{A}_{T}=\mathbf{0},\mathbf{b}_{T}=\begin{bmatrix}t_{x}\\ t_{y}\\ \end{bmatrix}$ (1)
where $\theta$ specifies the rotating angle. $t_{x}$ and $t_{y}$ indicate the vertical and horizontal translation distances (see Fig. 1).
Validation and Application Scope. Validating the above mutations is straightforward: developers can directly control the mutations via their parameters (e.g., $c$ in pixel-level mutations) and ensure generating valid test inputs by setting reasonable parameter values. These mutations are mostly adopted in testing safety critical VDL systems like autonomous driving [4, 5] due to their neat and explicit formulations. In addition, given these mutations’ applicability to almost all RGB images, they are widely used in testing more general VDL tasks like image classification [1].
@@ 3.2 Implicit & Data Exploration
Recent works diversify the mutated semantics with data-driven mutations. These mutations implicitly explore patterns in some images and then apply the patterns to mutate others.
@@
Style-level & Application Scope. From a high level, each image can be decomposed as its style and content, and VDL systems should primarily focus on the content rather than style. Thus, existing VDL testing works propose style-level mutations, which generate images of distinct styles but presumably retain the content of the seed image. The “style” has been studied from two aspects. First, the style can be viewed as describing how the image is produced, e.g., the seed image in Fig. 1 is taken in reality and its style is “natural”. In that sense, the style can be changed by generating an image’s artistic-stylized variants (e.g., an image having the same dog but painted using a Van Gogh style, as in Fig. 1). These mutations over “artistic styles” have been shown effective in flipping image classifier’s outputs in recent testing works  [37, 7], and some works also employ them to reveal VDL system’s bias to non-content semantics [38], showing their high value in VDL research.
Besides, some other works decompose the weather condition as the style and separate it from the main content. The weather conditions are transferred between images to generate test inputs, as shown in Fig. 1. However, these weather-targeted style-level mutations are specifically designed for testing VDL-based autonomous driving [6], and show limited applications in other testing scenarios.
Perceptual-level & Application Scope. Recent VDL testing focuses on fine-grained semantics of the objects in images and implements perceptual-level mutations [39, 40, 7], e.g., in Fig. 1, the dog’s eyes and mouth can be mutated to evaluate whether the VDL accurately recognizes different dogs [39]. Similar techniques can be applied to face photos for testing and improving finer-grained VDL tasks like face recognition [41, 42], face aging [43, 44], crowd counting [45, 46], etc.
Validation. Both style-level and perceptual-level mutations are implemented based on generative models like GANs [47, 6, 33]. To control the mutations, existing works propose optimization-based solutions, e.g., the cycle-consistency mechanism is formed during training in [47, 6] to better separate styles from images. However, such constraints derived from optimizations lack explicit guarantee, e.g., the cycle-consistency can be violated when transferring styles among images of dissimilar objects [48, 7]. Hence, existing testing works additionally adopt numeric metrics, such as Inception Score (IS) [49], Frechet Inception Distance (FID) [50], to measure the validity of generated test inputs. Since variations on fine-grained semantics in style- and perceptual-level mutations can hardly be captured via pixel values, recent works propose to measure the extents of achieved mutations via image embedding [^3] A representation of the image that captures its semantics. The term embedding may be alternatively dubbed as “feature” in some works. [51, 40, 7] (e.g., obtaining style and content embeddings, or using embedding models like CLIP [15, 51] to compare image semantics with text descriptions).